# Revolutionizing Large Language Model Training through Dynamic Parameter Adjustment

### v1

@@ 3 Theoretical analysis
4 Experiments
4.1 Experimental setup
We designed our experiments based on the settings described in [24] to benefit from established hyperparameter configurations. Our studies are carried out on the LLaMA model [31], with model sizes reduced to 130M, 250M, and 350M. The specific hyperparameters for these models are detailed in Table 1. We use Adam optimizer to train the model with $\beta_{1}=0.9,beta_{2}=0.999$. We use a cosine learning rate schedule with 100 warm-up steps and a total of 40,000 training steps. We employ the Adam optimizer for training, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The experiments utilize the C4 dataset [6], with the first 46M samples of the training dataset serving as our training data, and samples from the entire validation dataset used for testing. The evaluation of validation loss is performed on 10M tokens for all our experiments, with evaluations conducted every 1,000 steps. All experiments are conducted using 8xNVIDIA Tesla A100 80GB PCIe GPUs. Gradient accumulation is applied when GPU memory reaches its limit.
Table 1: Model sizes and architectures used in our experiments
Params Hidden Heads Layers Batch size Batch size per GPU Seq. len.
130M 768 12 12 600 300 256
250M 768 16 24 1152 72 512
350M 1024 16 24 1152 72 512
In our experiments involving LoRA and SwiLoRA, we set the LoRA rank to $128$ or $256$. This rank is chosen based on previous findings [38, 24, 39] suggesting it is an effective approximation for models smaller than 1B. To ensure fairness across all experiments, the initialization method described in Section 2 is applied to both LoRA and SwiLoRA experiments. We deploy LoRA adapters across all attention layers and fully connected layers in these experiments.
All experiments were repeated multiple times to select the best results. The learning rates were selected from a predefined set, including $0.0002,0.0005,\allowbreak 0.001,0.002,\allowbreak 0.005,0.01,\ldots$. The final chosen learning rates are detailed in Table 2.
Table 2: Learning rate settings for our experiments
130M 250M 350M
Full-rank 0.001 0.001 0.001
LoRA(rank$=128$) 0.01 0.01 0.01
SwiLoRA(rank$=128$) 0.02 0.02 0.02
[…]
4.2 Experimental results
Figures 2 displays the experimental results for the 130M, 250M, and 350M models, respectively, with the LoRA rank set to $128$. The data reveal that while LoRA alone does not yield satisfactory training results, SwiLoRA approaches the performance of full-rank training. The performance gap continues to grow as model size increases. This suggests that the low-rank training approach, such as LoRA, might cause models to become trapped in local minima, while SwiLoRA mitigates this issue by dynamically changing trainable parameters.
Table 3: Perplexity results at step 38,000 for 130M, 250M, and 350M models.
130M 250M 350M
Full-rank 27.71 20.19 18.72
LoRA(rank$=128$) 34.74 29.56 31.87
SwiLoRA(rank$=128$) 30.26 20.97 19.96
[…]
Figure 3: Loss results for 250M, and 350M models with a LoRA rank of $256$.
As shown in Figure 3, additional experiments conducted on the 250M and 350M models with a LoRA rank of $256$ demonstrates improved performance compared to those with the rank set at $128$, achieving outcomes close to those of full-rank training. Although utilizing a higher rank yields better outcomes, we argue that it is more economical to increase the model size for larger models for several reasons. Firstly, the method still has potential for further refinement. Secondly, a lower LoRA rank enables training on devices with limited memory capacities. Furthermore, in the context of 3D parallelism, inter-node communication is predominantly influenced by data parallelism, where communication overhead are proportional to trainable parameters. The trainable parameters for each model are detailed in Table 4. For further discussions on potential ways to enhance the SwiLoRA strategy, refer to Section 5. And the impact of distributed training is detailed in Appendix F.
Table 4: Comparison of trainable parameters: full-rank models vs. LoRA and SwiLoRA.
130M 250M/128 rank 350M/128 rank 250M/256 rank 350M/256 rank
Full-rank 134.3M 247.5M 368.2M 247.5M 368.2M
(Swi)LoRA 71.7M 98.9M 125.6M 148.4M 185.4M
In our experiments, we observed that switching frequencies that are either too large or too small, as well as the specific shape of the switching frequency curve, negatively affect training accuracy. The freezing steps constant $N$ also impacts the training accuracy. However, selecting an optimal value for $N$ is relatively straightforward, as this value is robust across different model since it simply determines how many steps are needed to warm up switched LoRA vectors. Detailed investigations into the effects of these hyperparameters can be found in Appendix C.
5 Limitations and future work
We hypothesize that the LoRA ranks specified in [24] are sufficient for training to achieve the same loss as full-rank training when the trainable parameters are frequently changed. However, experiments have demonstrated that the accuracy does not match that of full-rank training if the LoRA rank is not large enough. Besides, tuning the switching frequency of the LoRA vectors’ curve is challenging. To address these limitations, we propose the following directions for future work, as illustrated in Figure 4.
  • In our experiments, we simply used exponentially decreasing switching frequencies, which may not be the optimal approach. Conducting extensive experiments to determine the best frequencies across all steps is nearly impossible. Therefore, guidelines should be developed to help set appropriate switching frequencies throughout the training process.
  • Going further, a more detailed idea is to examine each layer of the model to adjust the switching frequencies. For instance, LoRA-drop [40] evaluates whether the rank is sufficient using a norm of $\Delta\mathbf{W}\mathbf{x}$. This is rational because different types of layers, such as the $Q,K,V$ matrices in transformer layers, exhibit significantly varied behaviors. Similarly, shallow and deep layers respond differently.
  • In our work, we simply chose candidate vectors at random. However, during training, all candidates are updated separately, leading to significant difference among them. The selection of these candidates may influence the training outcomes. Thanks to the relatively long switching interval(exceeding $40$ steps in our experiments), it is feasible to perform some time-consuming operations to determine the weights of the candidates for selection.
Figure 4: Future work roadmap.
6 Conclusions
In this work, we introduce SwiLoRA, a novel training strategy designed for parameter-efficient pre-training. Our approach achieves comparable accuracy to full-rank training while reducing the trainable parameters to approximately 50% to 60% of those in traditional full-rank training throughout the entire process. Moreover, the computational overhead and memory usage are nearly identical to those of LoRA when using the same number of trainable parameters. Additionally, we provide a theoretical analysis of our algorithm to confirm its correctness and to offer insights into its training methodology. The empirical results and theoretical analysis support our perspective on training: One can keep model parameters unchanged while change the part of trainable parameters during the training dynamically to achieve efficacy similar to full training.
Appendix A Related work
Similar to our approach, various LoRA variants employ strategies to increase the rank of a matrix by integrating weights into $\mathbf{W}$. For instance, ReLoRA [24] merges $\mathbf{B}\mathbf{A}$ into $\mathbf{W}$ and then restarts training at regular intervals. However, resetting all LoRA adapter parameters can lead to unstable training. To address this, ReLoRA introduces random optimizer pruning and a jagged schedule, but it still requires 33% of the steps to be full-rank training. Delta-LoRA [41], another variant, focuses on the fine-tuning phase. It updates the matrix $\mathbf{W}$ by using the gradients of the LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ simultaneously as they are updated. It improve accuracy for fine-tuning. However, it appears to be insufficient in increasing the rank of the modified parameters.
An attractive feature of LoRA is its support for quantization without loss of accuracy. This is achieved by training the LoRA adapter matrices $\mathbf{B}\mathbf{A}$ with high precision, while keeping the frozen part $\mathbf{W}$ at low precision within transformer layers. Examples of such quantization include QLoRA [5], LoftQ [21], and L4Q [17]. Additionally, the integration of quantization into SwiLoRA presents an intriguing topic. The feasibility of this integration is discussed in Appendix D.
On the other hand, LoRA and SwiLoRA are closely connected to SVD methods, although they do not require performing singular value decomposition. In fact, some studies utilize SVD to identify the most significant components of the decomposition of $\mathbf{B}\mathbf{A}$. For instance, earlier work by Pufferfish [34] and subsequent work in Cuttlefish[35] introduce adaptive strategies to determine the necessary steps for full-rank training and to select the rank of each linear layer for low-rank decomposition. InRank [38] develops a low-rank training approach based on greedy low-rank learning [22]. Another recent study, GaLore [39], employs SVD to apply a low-rank structure for approximating gradients, offering greater memory efficiency compared to LoRA. While these works are promising on pre-training, they require the computation of SVD, which has a time complexity of $O(n^{3})$ for $n\times n$ matrices. This computational demand becomes prohibitive for models with billions of parameters.
Table 5 provides a feature comparison among several related works discussed above.
Table 5: Comparison of different LoRA variations: LoRA, SwiLoRA, QLoRA and ReLoRA.
SwiLoRA LoRA QLoRA GaLore ReLoRA
Few (trainable) parameters ✓✓ ✓✓
Fast computation ✓✓ ✓✓ ✓✓ ✓✓
Easy implementation ✓✓
Effective at pre-training ✓✓ ✓✓
Appendix B Implementation of LoRA vector switching
The primary distinction in the implementation of SwiLoRA from conventional approaches lies in its handling of gradients and optimizer states at the granularity of row or column vectors within matrix parameters. Consider the scenario when using the AdamW optimizer: typically, each trainable parameter group in AdamW is associated with a “step” state which is implemented as a float scalar value in the code. To facilitate the resetting of specific rows or columns in matrices, we modify the type of "step" in the optimizer to a 32-bit float matrix with the same shape as the corresponding parameters. In fact, this modification does incur some extra memory overhead. An alternative approach would be to implement “step” as a row vector for $\mathbf{A}$ and a column vector for $\mathbf{B}$. However, this would require more complex code management, and thus, we have not adopted this strategy in our implementation.
With the capability to manipulate optimizer states and gradients at the level of rows and columns, we can now execute operations such as resetting optimizer states and freezing specific rows or columns of parameter matrices.
Additionally, note that all candidate vectors are stored in GPU memory, which consumes a modest amount of GPU resources. This portion of memory can be reduced by offloading the candidate vectors that are not actively in use to the CPU.
Table 6 presents a throughput comparison between SwiLoRA and LoRA under the same conditions described in Section 4, with the LoRA rank set to $128$. The results demonstrate that the throughput of SwiLoRA is comparable to that of LoRA in single GPU experiments. Additionally, experiments conducted on an 8-GPU setup reveal a more pronounced difference between LoRA and SwiLoRA, which may be attributed to CPU overhead. There is potential for further reductions in overhead through code optimization. This is feasible since all operations involve lower order time complexity.
Table 6: The throughput of SwiLoRA and LoRA for the 350M model, with a total batch size of 1152 and varying batch sizes per GPU for gradient accumulation. All experiments were conducted on a single A100 GPU, except for the final one, which utilized 8 A100 PCIe GPUs.
Batch sizes per GPU Batch size on 8 GPUs
1 18 36 72 72
LoRA 155s/step 16.5s/step 15.4s/step 14.8s/step 1.97s/step
SwiLoRA 159s/step 16.9s/step 15.9s/step 15.1s/step 2.37s/step
Appendix C Ablation study
@@
In Figure 6, we conduct a series of experiments with various frequency settings. The results indicate that the choice of frequency settings plays a crucial role in the model’s effectiveness. Specifically, we find that setting both the initial frequency values and the descent rates to moderate levels is essential for achieving optimal performance. Extremely high or low frequency settings tend to degrade the model’s performance, indicating a sensitive balance that must be maintained.
Figure 6: Perplexity comparison for the 130m model with different switching frequencies. The parameter $ratio$ determines the point at which the switching frequency is reduced to one-third of its initial value, occurring at the step $total\_step\times ratio$.
In Figure 7, we conduct experiments to investigate the impact of the number of frozen steps $N$. The results show that the choice of $N$ influences the loss outcomes. This can be explained as follows: when $N$ is excessively large, the training parameters may become biased towards different subsets of the data. Conversely, if $N$ is too small, at the moment the freezing is canceled, the gradients will have a larger contribution to the parameter updates due to the nature of momentum-based optimizers. This leads to potentially abrupt changes in model behavior.
@@
In Figure 8, we present the results from a focused comparative study where we evaluated our initialization strategy against the traditional LoRA initialization method through two distinct experiments. The results indicate that our initialization method outperform traditional approach for initialization. Notably, the loss curve for LoRA initialization reveals a slower decrease in initial loss compared to that of SwiLoRA initialization. This phenomenon in LoRA initialization can be attributed to the slow warm-up of matrix $\mathbf{A}$ and its associated candidate vectors due to (7). In contrast, our method modifies the initialization values to allow for more rapid adjustments, enabling the model to adapt more effectively to the training data.
@@
Appendix D Feasibility of quantization
Among all the LoRA variants that can be integrated into SwiLoRA, a quantization method such as QLoRA shows the most promise in enhancing efficiency and reducing memory usage. However, implementing quantization without significant accuracy loss is not straightforward. Given that LoRA vectors are frequently switched in SwiLoRA, the errors introduced by quantization can significantly impact the algorithm’s performance. For instance, when merging delta parameters triggered by the switching of LoRA vectors, the operation can be expressed as:
$\mathbf{W}\leftarrow\text{quant}(\text{dequant}(\mathbf{W})+\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime}).$ (18)
Here, precision errors may arise, with values approximately equal to
$\text{quant}(\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime{T}})-(\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime{T}}).$ (19)
Actually, it is possible to slightly adjust the values of candidate vectors before their selection to minimize this error. However, developing an efficient algorithm that effectively reduces this error without compromising accuracy is challenging. Therefore, we plan to explore this issue in future work.
@@ Appendix E Distribution of Singular values
Figure 10: Rank distribution of full-rank training and SwiLoRA on different types of linear layers.
Given that the rank distribution significantly influences the training efficacy of models [15, 7], we conducted experiments to examine the rank distribution of SwiLoRA. As outlined in Section 4, experiments were conducted on 350M model to analyze the rank distribution of linear layers after 40,000 training steps. Figure 9 demonstrates that the singular values of weight matrices converge within a limited range when trained with LoRA, indicating dominance of LoRA adapters in the linear layers. This dominance is expected, as the singular value distribution of weight matrices during the pre-training phase exhibits a form of illness, due to updates being limited to the low-rank adapter $\mathbf{B}\mathbf{A}$. In contrast, as illustrated in Figure 10, the rank distribution of SwiLoRA closely approximates that of full-rank training, suggesting a healthier and more effective adaptation process.

### v2

@@ Appendix A Theoretical analysis
Appendix B Ablation study
@@
In Figure 7, we conducted a series of experiments with various frequency settings. The results indicate that the choice of frequency settings plays a crucial role in the model’s effectiveness. Specifically, we find that setting both the initial frequency values and the descent rates to moderate levels is essential for achieving optimal performance. Extremely high or low frequency settings tend to degrade the model’s performance, indicating a sensitive balance that must be maintained.
Figure 7: Perplexity comparison for the 130m model with different switching frequencies. Each point in the figure has a triple label $(interval_{0},{ratio},{perplexity})$, with its size corresponding to the perplexity value. The parameter $ratio$ determines the point at which the switching frequency is reduced to one-third of its initial value, occurring at the step $total\_step\times ratio$.
In Figure 8, we conduct experiments to investigate the impact of the number of frozen steps $N$. The results indicate that the choice of $N$ influences the loss outcomes. This phenomenon can be explained as follows: when $N$ is excessively large, the training parameters may become biased towards different subsets of the data. Conversely, if $N$ is too small, at the moment the freezing is canceled, the gradients will have a larger contribution to the parameter updates due to the nature of momentum-based optimizers. This leads to potentially abrupt changes in model behavior. However, selecting an optimal value for $N$ is relatively straightforward, as this value is robust across different model since it simply determines how many steps are needed to warm up switched LoRA vectors. Therefore, this hyperparameter does not require frequent adjustments across various experiments.
@@
In Figure 9, we present the results from a focused comparative study where we evaluated our initialization strategy against the traditional LoRA initialization method through two distinct experiments. The results indicate that our initialization method outperform traditional approach for initialization. Notably, the loss curve for LoRA initialization reveals a slower decrease in initial loss compared to that of SwitchLoRA initialization. This phenomenon in LoRA initialization can be attributed to the slow warm-up of matrix $\mathbf{A}$ and its associated candidate vectors due to (6). In contrast, our method modifies the initialization values to allow for more rapid adjustments, enabling the model to adapt more effectively to the training data.
@@
Appendix C Experimental setting details
C.1 Experimental settings of ReLoRA
We adhere strictly to the setup described in ReLoRA [34] for our comparative experiments with ReLoRA. Specifically, the warm-up steps for the scheduler are set to 1,000. The learning rates are as follows: 5e-4 for full-rank pretraining, 1e-3 for ReLoRA, and 1e-2 for SwitchLoRA. The total batch size is established at 20,000. All other settings remain consistent with our previous experiments as detailed in Section 4.
C.2 Experimental settings of GaLore
Continuing in the same vein, we also strictly adhere to the setup outlined in GaLore [60] for our comparison experiments with GaLore. Specifically, we set the warm-up steps for the scheduler at 6,000. The total batch size is adjusted to 60,000. The learning rate is standardized at 1e-2 for all GaLore experiments, while for SwitchLoRA, it is set at 2e-2. All other experimental settings remain consistent with those detailed in our previous experiments, as described in Section 4.
C.3 Experimental settings of fine-tuning
The hyperparameters for fine-tuning used in the experiments described in Section 4.4 are presented in Table 8 and Table 9.
Table 8: Hyperparameters for different GLUE tasks for the 350M models.
CoLA STS-B MRPC RTE SST2
$lr$ (Full-rank) 8e-6 1e-5 1e-5 8e-6 3e-6
$lr$ (SwitchLoRA) 3e-5 5e-5 5e-5 5e-5 1e-5
$lr$ (GaLore) 2e-6 5e-6 5e-6 3e-6 8e-7
Batch size 16
Epochs 30
Sequence length 512
Table 9: Hyperparameters for different GLUE tasks for the 1.3B models.
CoLA STS-B MRPC RTE SST2
$lr$ (Full-rank) 8e-6 1e-5 2e-5 1e-5 5e-6
$lr$ (SwitchLoRA) 1e-5 1e-5 1e-5 2e-6 5e-6
Batch size 16
Epochs 30
Sequence length 512
Appendix D Implementation of LoRA vector switching
We discuss the code implementation of SwitchLoRA, focusing on its efficiency and memory consumption.
Implementation Adjustments in Optimizer
The primary distinction in the implementation of SwitchLoRA from conventional approaches lies in its handling of gradients and optimizer states at the granularity of row or column vectors within matrix parameters. Consider the scenario when using the AdamW optimizer: typically, each trainable parameter group in AdamW is associated with a “step” state which is implemented as a float scalar value in the code. To facilitate the resetting of specific rows or columns in matrices, we modify the type of "step" in the optimizer to a 32-bit float matrix with the same shape as the corresponding parameters. In fact, this modification does incur some extra memory overhead. An alternative approach would be to implement “step” as a row vector for $\mathbf{A}$ and a column vector for $\mathbf{B}$. However, this would require more complex code management, and thus, we have not adopted this strategy in our implementation. With the capability to manipulate optimizer states and gradients at the level of rows and columns, we can now execute operations such as resetting optimizer states and freezing specific rows or columns of parameter matrices.
Implementation of the Switching Process
We can either randomly select or sequentially select candidate vectors. However, fragmented operations on a GPU can’t fully utilize its capabilities. Since several candidate vectors are switched at each step, this will impact training efficiency. As an example, during the initial phase of SwitchLoRA training for the 1.3B LLaMA model with a LoRA rank of 512, approximately $\frac{512}{40}\approx 13$ candidate vectors are switched for each LoRA matrix at every step.
By organizing a list of candidate vectors into a matrix and selecting vectors sequentially, we can perform operations on multiple vectors simultaneously. For example, consider a scenario where we need to set the values of candidate vectors $\mathcal{C}(\mathbf{B})$ at indices 4, 5, 6 to the values of $\mathbf{B}$ at indices 7, 8, 9, respectively. Let $\mathcal{C}^{\mathbf{B}}$ be a matrix defined as $\mathcal{C}^{\mathbf{B}}\in\mathbb{R}^{m\times\min(m,n)}$, where each column $\mathcal{C}^{\mathbf{B}}{:,i}=\mathcal{C}(\mathbf{B})[i]$ for $i=1,\ldots,\min(m,n)$. We can then directly assign $\mathcal{C}^{\mathbf{B}}_{:,4:7}=\mathbf{B}_{:,7:10}$. This arrangement enables us to consolidate operations on multiple contiguous indices into a single operation, enhancing efficiency. Consequently, we employ a sequential selection approach and apply this technique. By implementing this approach, the switching process now occupies only about 1/40 of the training time during the initial training phase.
Memory offloading for candidate vectors
The use of candidate vectors leads to additional GPU memory usage. This memory overhead can be reduced by offloading it to the CPU. The offloading process can be decoupled from other training processes. By utilizing non-blocking CPU offloading, we can handle both offloading and other training processes in parallel, which can be readily implemented using frameworks like PyTorch.
The amount of parameters offloaded at each step is approximately $switch\_freq\times lora\_rank/hidden\_dim\times total\_param$. For the 1.3B LLaMA model, using 16-bit precision for model parameters, this translates to: $1/40\times 512/2048\times 1.3e9\times 2$ bytes $\approx 16.25MB$.
@@ Appendix E Distribution of Singular values
Figure 11: Rank distribution of full-rank training and SwitchLoRA on different types of linear layers.
Given that the rank distribution significantly influences the training efficacy of models [21, 11], we conducted experiments to examine the rank distribution of SwitchLoRA. As outlined in Section 4, experiments were conducted on the 350M model to analyze the rank distribution of linear layers after 40,000 training steps. Figure 10 demonstrates that the singular values of weight matrices converge within a limited range when trained with LoRA, indicating dominance of LoRA adapters in the linear layers. This dominance is expected, as the singular value distribution of weight matrices during the pre-training phase exhibits a form of illness, due to updates being limited to the low-rank adapter $\mathbf{B}\mathbf{A}$. In contrast, as illustrated in Figure 11, the rank distribution of SwitchLoRA closely approximates that of full-rank training, suggesting a more robust and more effective adaptation process.